You currently have $5,600. First United Bank will pay you an annual interest rate of 9.1, while Second National Bank will pay you an annual interest rate of 10.2. How many fewer years must you wait for your account value to grow to $15,700 at Second National Bank?

Multiple Choice

  • 1.22 years

  • 1.11 years

  • 1.47 years

  • 1.38 years

  • 1.13 years
Answer #1

Ans 1.22 years

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FIRST BANK
FV/ PV = (1 + r )^n
15700 / 5600 = (1 + 9.1%)^n
2.8035714285 = 1.091^n
n = 11.84 years
SECOND NATIONAL BANK
FV/ PV = (1 + r )^n
15700 / 5600 = (1 + 10.2%)^n
2.8035714285 = 1.102^n
n = 10.62 years
Difference = 11.84 - 10.62
1.22 years